Find, to the nearest tenth of a degree, all values of theta that satisfy the equation:
5 sine squared of theta-9 cosine of theta-3=0

Answer by lwsshak3(11628)

Find, to the nearest tenth of a degree, all values of theta that satisfy the equation:
5 sine squared of theta-9 cosine of theta-3=0
5sin^2(x)-9cosx-3=0
5(1-cos^2(x))-9cosx-3=0
5-5cos^2(x)-9cosx-3=0
5cos^2(x)+9cosx-2=0
(5cosx-1)(cosx+2)=0
..
cosx=2=0
cosx=-2
no solution: (-1 ≤ cosx ≤ 1)
or
5cosx-1=0
cosx=1/5
x=78.46˚+360˚n, 281.54˚+360˚n, n=any integer